OpenStudy (anonymous):

graph the parametric x(t)=1/3+12 y(t)=4t-3

OpenStudy (anonymous):

\[x(t)=\frac{t}{3}+12\]?

OpenStudy (anonymous):

if so solve for t an get \[x-12=\frac{t}{3}\] \[3x-36=t\] and then substitute this for t in the second expression. you will get y in terms of x, and have a line